OpenStudy (anonymous):

Will Give Medal Which function below has the lowest y intercept? f(x) graph going through (0, 5) and (2, 1) g(x) Kyle started the summer having read 3 books but plans to read 6 books per month over the summer. h(x) h(x) = 3x + 4 f(x) g(x) h(x) The functions all have the same y intercept.

OpenStudy (anonymous):

We know that linear equations have a slope (m) and y-intercept (b): y=mx+b, where x and y are variables. One point for f(x) is <0,5>, where x=0 and y=5. Another point is <4,3>. So we have two equations and two unknowns. We then can solve for m and b: 5=m(0)+b4=m(4)+b Using the 1st equation, we immediately see that b=5. This is the y=intercept for f(x). Make sense?

OpenStudy (anonymous):

No I Dont Get It Im Bad At This Can You Break It Down For Me Please

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

saying that she "started" the summer reading 3 books, means that at x=0, y=3, using our equation notation: 3=m(0) + b. This immediately shows the y-intercept for g(x).

OpenStudy (anonymous):

So its b = 3...

OpenStudy (anonymous):

So Is It B

OpenStudy (anonymous):

Yep ;D

OpenStudy (anonymous):

Thank You So Much

OpenStudy (anonymous):

Ay np

OpenStudy (anonymous):

Can YOu Help Me With Another Problem

OpenStudy (anonymous):

Sorry I gtg somewhere, I can help you when I'm online, tag me by doing @HydrgenCyanide

OpenStudy (anonymous):

Okay Thanks Again

OpenStudy (anonymous):

Np @audriim10